At a constant interest rate of 15%, compounded annually, what is the present value of an income stream paying $50 next quarter and growing at 2% per quarter until the end of the third year? From that point on it grows at 1% per quarter indefinitely.
Answer #1
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
Annual Interest rate=15%
Quarterly interest rate=i
(1+i)^4=1.15
1+i=1.15^(1/4)= 1.035558
i=0.035558
N=Quarter of Cash Flow
N A B=A/(1.03558^N)
Quarter Cash Flow Present Value
A 1 $50 $48.28
B=A*1.02 2 $51.00 $47.56
C=B*1.02 3 $52.02 $46.84
D=C*1.02 4 $53.06 $46.14
E=D*1.02 5 $54.12 $45.45
F=E*1.02 6 $55.20 $44.76
G=F*1.02 7 $56.31 $44.09
H=G*1.02 8 $57.43 $43.43
I=H*1.02 9 $58.58 $42.78
J=I*1.02 10 $59.75 $42.13
K=J*1.02 11 $60.95 $41.50
L=K*1.02 12 $62.17 $40.88 $533.84 SUM
M=L*1.01 13 $62.79 (62.17*1.01)
Value of income stream from quarter 13 to perpetuity at end of year 3 $2,456.78 (62.79/(0.035558-0.01)
Value of income stream from quarter 13 to perpetuity at PRESENT $1,615.37 2456.78/(1.15^3)
PRESENT VALUE OF TOTAL INCOME STREAM $2,149.21 (533.84+1615.37)